OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -3) and a directrix at y = 3.

OpenStudy (anonymous):

do you know what this looks like if so it is real easy

OpenStudy (anonymous):

i foget what it looks like

OpenStudy (anonymous):

lets draw a picture with just the focus and the directrix

OpenStudy (anonymous):



OpenStudy (anonymous):

the vertex is half way between the focus and the directrix, which is pretty clearly at the origin \((0,0)\)

OpenStudy (anonymous):

also from the fact that the focus is below the directrix, the parabola faces down

OpenStudy (anonymous):



OpenStudy (anonymous):

im following

OpenStudy (anonymous):

since the vertex is \((0,0)\) the equation will be \[4py=-x^2\] and all you need to finish is to find \(p\), the distance between the focus and the vertex

OpenStudy (anonymous):

should be more or less obvious that the distance is \(3\) so your equation is \[12y=-x^2\] or \[-12y=x^2\] or \[y=-\frac{x^2}{12}\] or whatever

OpenStudy (anonymous):

is the answer y^2= -12x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that is why it is important to know what it looks like before you start if it opens up or down, the \(x\) term is squared, not the \(y\) term

OpenStudy (anonymous):

my answer choices are: y2 = -12x y2 = -3x y = negative 1 divided by 12x2 y = negative 1 divided by 3x2

OpenStudy (anonymous):

did you see the answers i wrote above? one of them is there

OpenStudy (anonymous):

okay it's c y=-1/12 x^2

OpenStudy (anonymous):

as always, it is C

OpenStudy (anonymous):

haha thx again :)

OpenStudy (anonymous):

yw